OpenStudy (anonymous):

please write the equation of a line in general form! 1.Slope=-4/5 and y-intercept =1 2.slope =-8 and y-int =-5 3. through (1,-1) and slope=-5 PLEASE SHOW WORKINGS THANK YOU

OpenStudy (anonymous):

use the equation it is called the slope intercept form y=mx+c,where c=y intercept therefore 1.=5y+4x-5=0 2.y+8x+5=0 here the third one is different it is pont slope form (y-y1) = m(x-x1) therefore 3.y+5x-4=0

OpenStudy (anonymous):

in the last point x1and y1 are the points mentioned by u i.e(1,-1)